At a simple interest rate of 10% per year, determine how long will it take $10,000 to increase to 4 times as much. Compare the time it will take to double if the rate is 20% per year simple interest.

At an interest rate of 10%, it will take_____ years for the amount to increase 4 times as much.It will take______ years for the amount to double at an interest rate of 20%.
Answer #1

RATE OF INTEREST = 10%

PRINCIPAL = 10000

SO INTEREST EVERY YEAR = 10000 X 10% = 1000

NOW WE NEED 4 TIMES MORE THAN PRINCIPAL, SO INTEREST SHOULD BE 3 TIMES OF PRINCIPAL, WHICH IS EQUAL TO = 10000 X 3 = 30000

INTEREST PER YEAR = 1000

SO IT WILL TAKE 30 YEARS [30000/1000]

ANSWER 1 : 30 YEARS

NOW IF INTEREST RATE = 20%

INTEREST PER YEAR = 10000 X 20% = 2000

NOW WE NEED DOUBLE THE PRINCIPAL = 20000, SO INTEREST TO BE EARNED = 20000 -10000 = 10000

INTEREST PER YEAR = 2000

SO IT WILL TAKE 5 YEARS [10000/2000 = 5 YEARS]  TO EARN INTEREST OF 10000

ANSWER 2 : 5 YEARS [[Thumbs up please]